Grade 12Physical Chemistry

density of a 2.05M solution of acetic acid in water is 1.02 g/mL.The molality of the solution is??

sajid
9 Years agoGrade 12

1 Answer

Vikas TU
9 Years ago
               Let us find the weight of the acetic acid
   = 2.05 * 60
    =123
The weight of the solution
   =1000* 1.02
    =1020
         Weight of the water =(1020-123)
                                           =897g
               Therefore, Molarity is given as;
                        =2.05 *1000897
                        =2.285molkg-1